# Real enumerative invariants relative to the toric boundary and their refinement

### v1

@@ 5.2 Walls in the space of constraints
Thus, we are left with the case of $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$ consisting of $n$ distinct points and $\widehat{C}$ either consisting of two components, or being a rational curve with a node. Observe that the case of two rational components is not possible, since by Lemma 2.4, the dimension of such a stratum would not exceed $(n-2)+1=n-1<n$.
@@
Suppose that $\widehat{C}$ is as in item (2iv). For the dimension reason, the sequence $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$ is in general position subject to exactly two Menelaus type relations, while $w_{0}(t^{*})$ is in general position in ${\mathcal{Q}}$. Thus, by Lemma 2.4, we obtain the immersion and smoothness statements as required. The rest of the argument literally coincides with the corresponding part of the proof of Lemma 4.2(2iii).
@@
Suppose that $\widehat{C}$ is a rational curve with a node $p$. Then, this node cannot be mapped to ${\operatorname{Tor}}(P_{\Delta})^{\times}$, since otherwise one would encounter a real rational curve with two one-dimensional branches, one in ${\operatorname{Tor}}_{\mathbb{R}}^{+}(P_{\Delta})$ and the other in $\overline{\mathcal{Q}}$. Hence the node is mapped to $w\in\widehat{\boldsymbol{w}}_{\partial}(t^{*})$, and for the above reason, the intersection multiplicities of the two branches of $C={\Phi}(\widehat{C})$ at $w$ are odd. Observe that ${\Phi}^{\prime}:{\mathbb{P}}^{1}\to\widehat{C}\to{\operatorname{Tor}}(P_{\Delta})$ is not a multiple covering of the image. Indeed, in such a case, one would have that $P$ is a triangle, $l_{1}=l_{2}=l$, and $C={\Phi}(\widehat{C})$ is a rational curve intersecting at least two of the toric divisors with odd multiplicity, the double covering being ramified at the two points of $\widehat{\boldsymbol{w}}_{\partial}(t^{*})\setminus\{w\}$. With an automorphism of ${\mathbb{Z}}^{2}$, we can turn the Newton triangle of $C$ (which is $\frac{1}{2}P_{\Delta}$) into the triangle
@@
Since $x(1)>0$, we get $a>0$, and hence $x>0$ as $0<\theta\ll 1$. Then $y>0$ for $0<\theta\ll 1$, and hence $b<0$. Thus, if $r$ and $s$ are odd, we obtain $x>0,y<0$ as $\theta\to+\infty$, and $x<0,y<0$ as $\theta\to-\infty$. If $r$ is even and $s$ is odd, we get $x>0,y<0$ as $\theta\to+\infty$, and $x<0,y>0$ as $\theta\to-\infty$. In both cases, it follows that the intersection point with the toric divisor ${\operatorname{Tor}}([(0,s),(r+2l+1,0)])$ is out of $\partial{\operatorname{Tor}}_{\mathbb{R}}^{+}(P_{\Delta})$ against the initial assumption. If $r$ is odd and $s$ is even, we get that the segment of ${\mathbb{R}}C$ in the quadrant $\overline{\{x>0,y<0\}}$ touches all three toric divisors, which contradicts the condition (AQC). Thus, a multiple covering is excluded. The rest of claims can be derived in the same way as the statement of Lemma 4.2(2ii) with one exception: we shall show that the local branches of $C$ at $w$ intersect each other with multiplicity $\min\{2l_{1}+1,2l_{2}+1\}$. For $l_{1}\neq l_{2}$, this is immediate. Suppose that $l_{1}=l_{2}=l$. If $n=3$, then $P$ is a triangle and, in the above setting, we get the parametrization of $C$ in the form
@@
Suppose now that $\widehat{C}$ is a smooth elliptic curve and ${\Phi}:\widehat{C}\to{\operatorname{Tor}}(P_{\Delta})$ is a multiple covering of its image. Since the preimages of at least $n-2$ points of $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$ in $\widehat{C}$ are irreducible, the map ${\Phi}:\widehat{C}\to C\subset{\operatorname{Tor}}(P_{\Delta})$ is ramified, and hence $C$ cannot be elliptic by the Riemann-Hurwitz formula. Thus, $C$ is rational, which for dimension reason implies that $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$ consists of $n$ points in general position subject to one Menelaus condition. In particular, $C$ is smooth at each point of $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$, and ${\Phi}$ has ramification index $s$ at all points of $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$. By the Riemann-Hurwitz formula,
@@
[…]
@@ 5.3 Wall-crossing
Let $\{(\widehat{\boldsymbol{w}}_{\partial}(t),w_{0}(t))\}_{0\leq t\leq 1}$ be a generic path in $\operatorname{Men}_{\mathbb{R}}^{\tau}\times{\mathcal{Q}}$, and let $t^{*}\in(0,1)$ be such that ¯$\overline{\mathcal{M}}_{1,n}^{\mathbb{R}}(\Delta,\widehat{\boldsymbol{w}}(t^{*}))$ contains an element $\xi$ as described in one of the items of Lemma 5.2(2). Then, for each $\kappa\in\frac{1}{2}{\mathbb{Z}}$ such that $|\kappa|\leq{\mathcal{A}}(\Delta)$, the numbers $W_{1}^{\kappa}(t):=W_{1}^{\kappa}(\Delta,\widehat{\boldsymbol{w}}(t))$ do not change as $t$ varies in a neighborhood of $t^{*}$.
@@
(1) Suppose that $\xi$ is as in Lemma 5.2(2i). Consider the lifts of $w_{0}(t^{*})$ upon ${\mathbb{E}}$ that correspond to real branches of $C$ centered at $w_{0}(t^{*})$. Since $C$ is immersed, we only have to show that each of the lifts induces a germ of a smooth one-dimensional subfamily in ${\mathcal{M}}_{1,n}({\operatorname{Tor}}(P_{\Delta}),{\mathcal{L}}_{P_{\Delta}})$. This follows from the smoothness statements in [3, Proposition 4.17(2)] and in [24, Lemma 3(1)], [13, Lemma 2.4(2)].
(2) Suppose that $\xi$ is as in Lemma 5.2(2ii). The argument used in step (1) of the proof of Lemma 4.3 applies, provided we establish a relevant transversality statement. Here, the transversality condition is similar to (12):
$h^{0}({\mathbb{E}},{\mathcal{O}}_{\mathbb{E}}({\boldsymbol{d}}-p_{0}))=0,\quad{\boldsymbol{d}}\in{\operatorname{Div}}({\mathbb{E}})\ ,$ (23)
@@
$=C^{2}-(C^{2}-c_{1}({\operatorname{Tor}}(P_{\Delta}))c_{1}({\mathcal{L}}_{P_{\Delta}}))-(c_{1}({\operatorname{Tor}}(P_{\Delta}))c_{1}({\mathcal{L}}_{P_{\Delta}})-1)=1>0\ .$
Observe that by the Riemann-Roch formula, $h^{0}({\mathbb{E}},{\mathcal{O}}_{\mathbb{E}}({\boldsymbol{d}}))=1$, and hence (23) holds due to the general position of $w_{0}$ on $C$.
(3) The treatment of the case where $\xi$ is as in Lemma 5.2(2iii) literally coincides with that in step (2) of the proof of Lemma 4.3, when we replace the transversality relation (12) with (23).
(4) Suppose that $\xi$ is as in Lemma 5.2(2iv). Due to the condition (AQC), the real part of the rational curve $C_{1}={\Phi}(\widehat{C}_{1})$ must lie in ${\operatorname{Tor}}_{\mathbb{R}}^{+}(P_{\Delta})$, and the real part of the elliptic curve $C_{2}={\Phi}(\widehat{C}_{2})$ has two one-dimensional components, one in ${\operatorname{Tor}}_{\mathbb{R}}^{+}(P_{\Delta})$ and the other in $\overline{\mathcal{Q}}$. Then, the local argument proving the constancy of the numbers $W_{1}^{\kappa}(t)$ literally coincides with that in step (5a) of the proof of Lemma 4.3.
The global argument is similar to that in step (5b) of the proof of Lemma 4.3. Here, we obtain the conditions
[…]
and
[…]
The former one follows by the Riemann-Roch formula, and the latter one as in part (2) above.
@@
(5) Suppose that $\xi$ is as in Lemma 5.2(2v). Since $\widehat{C}$ is rational, both local branches $B_{1},B_{2}$ of $C={\Phi}(\widehat{C})$ at $w=w_{i}^{\sigma}(t^{*})$ are real and each one intersects ${\operatorname{Tor}}(\sigma)$ with odd multiplicities $12l_{1}+1,2l_{2}+$ 1, respectively, where $2l_{1}+2l_{2}+2=2k_{i}^{\sigma}$, $l_{1}\leq l_{2}$. Choose two mobile points $w_{i^{\prime}}^{\sigma^{\prime}},w_{i^{\prime\prime}}^{\sigma^{\prime\prime}}\in\widehat{\boldsymbol{w}}_{\partial}(t^{*})\setminus\{w\}$. The one-dimensional stratum {¯$\{\overline{\mathcal{M}}_{1,n}^{{\mathbb{R}},+}(\Delta,{\boldsymbol{w}}(t))\}_{|t-t^{*}|<\zeta}$, $0<\zeta\ll 1$, projects to the germ at $C$ of a one-dimensional variety in $|{\mathcal{L}}_{P_{\Delta}}|$, and denote by ${\mathcal{V}}$ one of its real irreducible components. Let the tangent line to ${\mathcal{V}}$ at $C$ be spanned by $C$ and $C^{*}\in|{\mathcal{L}}_{P_{\Delta}}|\setminus\{C\}$. Fix equavariant coordinates $x,y$ in a neighborhood of $w$ so that $w=(0,0)$ and ${\operatorname{Tor}}(\sigma)=\{y=0\}$. Then, $C$ in a neighborhood of $w$ is given by an equation
@@
$y^{2}+P(x)y+\eta_{1}\eta_{2}x^{4l+2}+\text{h.o.t.}=0,\quad P(x)=(\eta_{1}+\eta_{2})x^{2l+1}+\sum_{j=2l}^{0}a_{j}t_{1}^{2l+1-j}x^{j}\ ,$ (24)
$\text{where}\quad t_{1}=t-t^{*},\quad\text{and}\quad x^{2l+1}P(1/x)=\lambda_{1}{\operatorname{Cheb}}_{2l+1}(t_{1}x+\lambda_{2})+\lambda_{3}$
with real numbers $\lambda_{1},\lambda_{2},\lambda_{3}$ uniquely determined by $\eta_{1},\eta_{2}$ and the coefficient of $x^{2l}y$ in the equation of $C^{*}$, and h.o.t. includes higher powers of $t_{1}$ in the present monomials and the sum of monomials above ${\Gamma}$. If $\eta_{1}\eta_{2}>0$, the latter formula yields $2l$ elliptic nodes in a neighborhood of $w$. Observe that the change of sign of $t_{1}$ yields the symmetry with respect to the origin for the real part of $C^{(t)}$ (up to h.o.t.), and hence (see Figure 1(c)) the quantum index and the Welschinger sign (9) persists along the path $\{(\widehat{\boldsymbol{w}}_{\partial}(t),w_{0}(t))\}_{|t-t^{*}|<\zeta}$. If $\eta_{1}\eta_{2}<0$, then equation (24) implies
@@
which means that $C^{(t)}$ has only non-real nodes in a neighborhood of $w$. Again (see Figure 1(d)), the quantum index and the Welschinger sign remain unchanged.
@@
At last, to have the family ${\mathcal{V}}$ smooth, representing the local deformation as above, and crossing the wall transversally, we need the following (global) transvrsality condition: the Zariski tangent space to the germ at $C^{*}$ of the family constrained by the given tangency conditions at fixed points of $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$ and at two mobile points of $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$ as well as by fixing the point $w_{0}$ has dimension zero. The routine computations, performed in preceding steps, reduce this condition to

### v2

@@ 5.2 Walls in the space of constraints
(2iv) We are left with the case of $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$ consisting of $n$ distinct points and $\widehat{C}$ either consisting of two components, or being a rational curve with a node. Observe that the case of two rational components is not possible, since a generic element of ${\mathcal{M}}^{\mathbb{R}}_{0,n}(\Delta,\widehat{\boldsymbol{w}}(t))$ will appear after smoothing out two intersection points of the given rational curves, and hence the pairs of rational curves would sweep a locus of codimension $2$.
@@
Now suppose that $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$ consists of $n$ distinct points, $\widehat{C}=\widehat{C}_{1}\cup\widehat{C}_{2}$, where $\widehat{C}_{1}\simeq{\mathbb{P}}^{1}$ and $\widehat{C}_{2}$ is a smooth elliptic curve, and the intersection $\widehat{C}_{1}\cap\widehat{C}_{2}$ consists of one point $p$. Then, for the dimension reason, the sequence $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$ is in general position subject to exactly two Menelaus type relations, while $w_{0}(t^{*})$ is in general position in ${\mathcal{Q}}$. Thus, by Lemma 2.4, we obtain the immersion and smoothness statements as required. The rest of the argument literally coincides with the corresponding part of the proof of Lemma 4.2(2iii).
@@
(2v) Suppose that $\widehat{C}$ is a rational curve with a node $p$. Then, this node cannot be mapped to ${\operatorname{Tor}}(P_{\Delta})^{\times}$, since otherwise one would encounter a real rational curve with two one-dimensional branches, one in ${\operatorname{Tor}}_{\mathbb{R}}^{+}(P_{\Delta})$ and the other in $\overline{\mathcal{Q}}$. Hence the node is mapped to $w\in\widehat{\boldsymbol{w}}_{\partial}(t^{*})$, and for the above reason, the intersection multiplicities of the two branches of $C={\Phi}(\widehat{C})$ at $w$ are odd. Observe that ${\Phi}^{\prime}:{\mathbb{P}}^{1}\to\widehat{C}\to{\operatorname{Tor}}(P_{\Delta})$ is not a multiple covering of the image. Indeed, in such a case, one would have that $P$ is a triangle, $l_{1}=l_{2}=l$, and $C={\Phi}(\widehat{C})$ is a rational curve intersecting at least two of the toric divisors with odd multiplicity, while the covering is double ramified at the two points of $\widehat{\boldsymbol{w}}_{\partial}(t^{*})\setminus\{w\}$. With an automorphism of ${\mathbb{Z}}^{2}$, we can turn the Newton triangle of $C$ (which is $\frac{1}{2}P_{\Delta}$) into the triangle
@@
Since $x(1)>0$, we get $a>0$, and hence $x>0$ as $0<\theta\ll 1$. Then $y>0$ for $0<\theta\ll 1$, and hence $b<0$. Thus, if $r$ and $s$ are odd, we obtain $x>0,y<0$ as $\theta\to+\infty$, and $x<0,y<0$ as $\theta\to-\infty$. If $r$ is even and $s$ is odd, we get $x>0,y<0$ as $\theta\to+\infty$, and $x<0,y>0$ as $\theta\to-\infty$. In both cases, it follows that the intersection point with the toric divisor ${\operatorname{Tor}}([(0,s),(r+2l+1,0)])$ is out of $\partial{\operatorname{Tor}}_{\mathbb{R}}^{+}(P_{\Delta})$ against the initial assumption. If $r$ is odd and $s$ is even, we get that the segment of ${\mathbb{R}}C$ in the quadrant $\overline{\{x>0,y<0\}}$ touches all three toric divisors, which contradicts the condition (AQC). Thus, a multiple covering is excluded. The rest of claims can be derived in the same way as the statement of Lemma 4.2(2ii) with one exception: we show that the local branches of $C$ at $w$ intersect each other with multiplicity $\min\{2l_{1}+1,2l_{2}+1\}$. For $l_{1}\neq l_{2}$, this is immediate. Suppose that $l_{1}=l_{2}=l$. If $n=3$, then $P$ is a triangle and, in the above setting, we get the parametrization of $C$ in the form
@@
(2vi) Suppose now that $\widehat{C}$ is a smooth elliptic curve and ${\Phi}:\widehat{C}\to{\operatorname{Tor}}(P_{\Delta})$ is a multiple covering of its image. Since the preimages of at least $n-2$ points of $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$ in $\widehat{C}$ are irreducible, the map ${\Phi}:\widehat{C}\to C\subset{\operatorname{Tor}}(P_{\Delta})$ is ramified, and hence $C$ cannot be elliptic by the Riemann-Hurwitz formula. Thus, $C$ is rational, which for dimension reason implies that $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$ consists of $n$ points in general position subject to one Menelaus condition. In particular, $C$ is smooth at each point of $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$, and ${\Phi}$ has ramification index $s$ at all points of $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$. By the Riemann-Hurwitz formula,
@@
Hence, $n\leq\frac{2s}{s-1}$,which in view of the assumption $n\geq 4$, yields $s=2$ and $n=4$ as stated in item (vi). $\Box$
@@ 5.3 Wall-crossing
Let $\{(\widehat{\boldsymbol{w}}_{\partial}(t),w_{0}(t))\}_{0\leq t\leq 1}$ be a generic path in $\operatorname{Men}_{\mathbb{R}}^{\rho}\times{\mathcal{Q}}$, and let $t^{*}\in(0,1)$ be such that ¯$\overline{\mathcal{M}}_{1,n}^{\mathbb{R}}(\Delta,\widehat{\boldsymbol{w}}(t^{*}))$ contains an element $\xi$ as described in one of the items of Lemma 5.2(2). Then, for each $\kappa\in\frac{1}{2}{\mathbb{Z}}$ such that $|\kappa|\leq{\mathcal{A}}(\Delta)$, the numbers $W_{1}^{\kappa}(t):=W_{1}^{\kappa}(\Delta,\widehat{\boldsymbol{w}}(t))$ and $\widetilde{W}_{1}^{\kappa}(t):=\widetilde{W}_{1}^{\kappa}(\Delta,\widehat{\boldsymbol{w}}(t))$ do not change as $t$ varies in a neighborhood of $t^{*}$.
@@
(1) Suppose that $\xi$ is as in Lemma 5.2(2i). Consider the lifts of $w_{0}(t^{*})$ upon $\widehat{C}$ that correspond to real branches of $C$ centered at $w_{0}(t^{*})$. Since $C$ is immersed, we only have to show that each of the lifts induces a germ of a smooth one-dimensional subfamily in ${\mathcal{M}}_{1,n}({\operatorname{Tor}}(P_{\Delta}),{\mathcal{L}}_{P_{\Delta}})$. This follows from the smoothness statements in [4, Proposition 4.17(2)] and in [30, Lemma 3(1)], [16, Lemma 2.4(2)].
(2) Suppose that $\xi$ is as in Lemma 5.2(2ii). We follow the argument used in Steps (1) and (2) of the proof of Lemma 4.3. Since the local treatment of the occurring bifurcations is the same as in the proof of Lemma 4.3, we derive the constancy of $W_{1}^{\kappa}(t)$ and $\widetilde{W}_{1}^{\kappa}(t)$ along the interval $(t^{*}-{\varepsilon},t^{*}+{\varepsilon})$ from the fact that the image of $H^{0}(C,{\operatorname{in}}*{\mathcal{L}}_{P_{\Delta}})$ in $\prod_{\sigma,i}\mathrm{Vers}(C,w_{i}^{\sigma})\times\prod_{z\in{\operatorname{Sing}}\hskip 1.70709pt(C)\cap{\operatorname{Tor}}(P_{\Delta})^{\times}}\mathrm{Vers}(C,z)$ transversally intersects with the tangent cone to the locus (19), where $w_{i}^{\sigma}$ is a center of a local branch of $C$ of type $A_{2m}$, $m\geq k_{i}^{\sigma}$. This amounts to the relation of the form (20) which we rewrite as
$h^{1}(\widehat{C},{\mathcal{O}}_{\widehat{C}}({\boldsymbol{d}}_{1}))=0,$ (36)
where the divisor ${\boldsymbol{d}}_{1}$ satisfies
$\deg{\boldsymbol{d}}_{1}=C^{2}-(C^{2}+CK_{{\operatorname{Tor}}(P_{\Delta})})-(-CK_{{\operatorname{Tor}}(P_{\Delta})}-n+1)-({\operatorname{mult}}(C,w_{i}^{\sigma})-1)=n-2>0,$
and ${\operatorname{mult}}(C,w_{i}^{\sigma})$ stands for the multiplicity of $C$ at $w_{i}^{\sigma}$. Thus, the required transversality follows.
(3) The treatment of the case where $\xi$ is as in Lemma 5.2(2iii) is the same as that in step (3) of the proof of Lemma 4.3. The required transvesality condition reads now as (cf. (21))
$h^{1}(\widehat{C},{\mathcal{O}}_{\widehat{C}}({\boldsymbol{d}}))=0,$
@@
$\deg{\boldsymbol{d}}=C^{2}-((C^{2}+CK_{{\operatorname{Tor}}(P_{\Delta})})-(-CK_{{\operatorname{Tor}}(P_{\Delta})}-2)=2>0,$
which immediately yields the above $h^{1}$-vanishing.
(4) Suppose that $\xi$ is as in Lemma 5.2(2iv) and ${\Phi}(p)\in{\operatorname{Tor}}(P_{\Delta})^{\times}$. The real part of the elliptic curve $C_{2}={\Phi}(\widehat{C}_{2})$ has two one-dimensional components, one in ${\operatorname{Tor}}_{\mathbb{R}}^{+}(P_{\Delta})$ and the other in $\overline{\mathcal{Q}}$. The real part of the rational curve $C_{1}$ lies either in ${\operatorname{Tor}}_{\mathbb{R}}^{+}(P_{\Delta})$, or in $\overline{\mathcal{Q}}$. Again the treatment of the wall-crossing is very similar to that in step (4) of the proof of Lemma 4.3. We only comment on the transversality conditions ensuring the realizability of this scenario of the wall-crossing ina suitable variation of the constraints. Thus, we have a combination of two $h^{1}$-vanishinhg relations: (i) one is for $\widehat{C}_{1}$, and it coincides with (22), (ii) the other reads (cf. (22))
$h^{1}(\widehat{C}_{2},{\mathcal{O}}_{C_{2}}({\boldsymbol{d}}_{2}))=0,$
where
$\deg{\boldsymbol{d}}_{2}=C_{2}^{2}-(C_{2}^{2}+CK_{{\operatorname{Tor}}(P_{\Delta})})-(-CK_{{\operatorname{Tor}}(P_{\Delta})}-1)=1>0,$
which yields the desired $h^{1}$-vanishing.
Suppose that $\xi$ is as in Lemma 5.2(2iv) and ${\Phi}(p)=w_{i}^{\sigma}\in\widehat{\boldsymbol{w}}(t^{*})$ (to relax notation, below we write $w$ instead of $w^{\sigma}_{i}$). Again, the local description of the corresponding wall-crossing event can be borrowed from part (5) of the proof of Lemma 4.3 and correspondingly, the signs $W_{1}^{\kappa}(\widehat{\boldsymbol{w}}(t))$ and $\widetilde{W}_{1}^{\kappa}(\widehat{\boldsymbol{w}}(t))$ do not change as $t$ transints through $t^{*}$. Thus, we should confirm the required transversality conditions.
If ${\Phi}:\widehat{C}_{1}\to C_{1}$ and ${\Phi}:\widehat{C}_{2}\to C_{2}$ both are immersions, then we need the relation (24) which reduces to two $h^{1}$-vanishong conditions (25) on $\widehat{C}_{1}\simeq{\mathbb{P}}^{1}$ and $\widehat{C}_{2}$, a smooth elliptic curve. For $\widehat{C}_{1}$, that $h^{1}$-vanishing holds due to (26). For $\widehat{C}_{2}$, it reads
$h^{1}(\widehat{C}_{2},{\mathcal{O}}_{C_{2}}({\boldsymbol{d}}_{2}))=0,$ (37)
where the divisor ${\boldsymbol{d}}_{2}$ satisfies
$\deg{\boldsymbol{d}}_{2}=C_{2}^{2}-(C_{2}^{2}+C_{2}K_{{\operatorname{Tor}}(P_{\Delta})})-(-C_{2}K_{{\operatorname{Tor}}(P_{\Delta})}-|C_{2}\cap{\operatorname{Tor}}(\partial P_{\Delta}|+1)$
$=|C_{2}\cap{\operatorname{Tor}}(\partial P_{\Delta}|-1\geq 2>0,$
and hence (37) holds too.
In the case of ${\Phi}:\widehat{C}_{1}\to C_{1}$ a $2l_{1}$-multiple ramified covering and ${\Phi}:\widehat{C}_{2}\to C_{2}$ an immersion, the required transversality condition takes the form of (27). Using the Horace method in the same manner as in part (5b) of the proof of Lemma 4.3, we come to the system of $h^{1}$-vanishing relations (28). The upper line relations in (28) are the same in the current setting, and hence hold for the same reason, while the second line relation turns into
$h^{1}(\widehat{C}_{2},{\mathcal{O}}_{\widehat{C}_{2}}({\boldsymbol{d}}))=0,$
where
$\deg{\boldsymbol{d}}=C_{2}^{2}-(C_{2}^{2}+C_{2}K_{{\operatorname{Tor}}(P_{\Delta})})-(-C_{2}K_{{\operatorname{Tor}}(P_{\Delta})}-1)=1>0,$
and hence it holds too.
@@
(5) Suppose that $\xi$ is as in Lemma 5.2(2v). Since $\widehat{C}$ is rational, both local branches $B_{1},B_{2}$ of $C={\Phi}(\widehat{C})$ at $w=w_{i}^{\sigma}(t^{*})$ are real and they intersect ${\operatorname{Tor}}(\sigma)$ with odd multiplicities $12l_{1}+1,2l_{2}+$ 1, respectively, where $2l_{1}+2l_{2}+2=2k_{i}^{\sigma}$, $l_{1}\leq l_{2}$. Choose two mobile points $w_{i^{\prime}}^{\sigma^{\prime}},w_{i^{\prime\prime}}^{\sigma^{\prime\prime}}\in\widehat{\boldsymbol{w}}_{\partial}(t^{*})\setminus\{w\}$ and fix the other points of $\widehat{\boldsymbol{w}}(t^{*})$. The one-dimensional stratum {¯$\{\overline{\mathcal{M}}_{1,n}^{{\mathbb{R}},+}(\Delta,{\boldsymbol{w}}(t))\}_{|t-t^{*}|<\zeta}$, $0<\zeta\ll 1$, projects to the germ at $C$ of a one-dimensional variety in $|{\mathcal{L}}_{P_{\Delta}}|$, and we denote by ${\mathcal{V}}$ one of its real irreducible components. Let the tangent line to ${\mathcal{V}}$ at $C$ be spanned by $C$ and $C^{*}\in|{\mathcal{L}}_{P_{\Delta}}|\setminus\{C\}$. Fix conjugation-invariant coordinates $x,y$ in a neighborhood of $w$ so that $w=(0,0)$ and ${\operatorname{Tor}}(\sigma)=\{y=0\}$. Then, in a neighborhood of $w$, the curve $C$ is given by an equation
@@
$y^{2}+P(x)y+\eta_{1}\eta_{2}x^{4l+2}+\text{h.o.t.}=0,\quad P(x)=(\eta_{1}+\eta_{2})x^{2l+1}+\sum_{j=2l}^{0}a_{j}\tau^{2l+1-j}x^{j}\ ,$ (38)
$\text{where}\quad\tau=t-t^{*},\quad\text{and}\quad x^{2l+1}P(1/x)=\lambda_{1}{\operatorname{Cheb}}_{2l+1}(\tau x+\lambda_{2})+\lambda_{3}$
with real numbers $\lambda_{1},\lambda_{2},\lambda_{3}$ uniquely determined by $\eta_{1},\eta_{2}$ and by the coefficient of $x^{2l}y$ in the equation of $C^{*}$, and h.o.t. includes higher powers of $\tau$ in the present monomials and the sum of monomials above $\Gamma$. If $\eta_{1}\eta_{2}>0$, the latter formula yields $2l$ elliptic nodes in a neighborhood of $w$, all of them lying either in the domain $\{y>0\}$, or in the domain $\{y<0\}$. Observe that the change of sign of $\tau$ yields the symmetry with respect to the origin for the real part of $C^{(t)}$ (up to h.o.t.), while the orientation of the local branch of $C^{(t)}$ at $w$ persists, and hence the quantum index and the signs $W_{1}^{\kappa}(t)$ and $\widetilde{W}_{1}^{\kappa}(t)$ remain constant along the path $\{(\widehat{\boldsymbol{w}}(t))\}_{|t-t^{*}|<\zeta}$. If $\eta_{1}\eta_{2}<0$, then equation (38) implies
@@
which means that $C^{(t)}$ has only non-real nodes in a neighborhood of $w$. Again, the orientation of the local branch of $C^{(t)}$ at $w$ persists (see Figure 6(d)), and we conclude that the signs $W_{1}^{\kappa}(t)$ and $\widetilde{W}_{1}^{\kappa}(t)$ do not change as $t$ transits through $t^{*}$.
@@
At last, we establish the transversality condition sufficient for the realizability of the wall-crossing described above by a variation of constraints. Namely, we prove that representing the local deformation as above, and crossing the wall transversally, we need the following transversality condition: the Zariski tangent space to the germ at $C^{*}$ of the family constrained by the given tangency conditions at fixed points of $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$ and at two mobile points of $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$ as well as by fixing the point $w_{0}$ has dimension zero. The routine computations, performed in preceding steps, reduce this condition to